# Quantum critical theories in a periodic potential: strange metallic thermoelectric and magneto-transport

### v1

@@ 5 Thermoelectric transport in 2D charge neutral lattices
We will show that there is a qualitatively novel contribution to the phenomenology of quantum critical transport in 2D charge-neutral lattices. In addition the electrical conductivity has a modified RG-scaling from 1D. At the same time, the novel pole-collision physics behind 1D thermal transport does appear to continue to hold in 2D.
@@ 5.1 DC transport in 2D charge-neutral lattices
The difference in electrical transport is immediately evident in the DC conductivity $\sigma_{DC}$, as shown in Figure 6 (panel (a)): the electrical conductivity now increases with temperature instead of decreasing monotonically back down to the translationally invariant quantum critical value $\sigma_{\infty}$. To explain this remarkable enhanced conductivity, one can make an analogy between a periodic source of translational symmetry breaking which vanishes averaged over a lattice cell $\int_{\mathbf{x}}^{\mathbf{x}+\mathbf{a}}\mu(\mathbf{x})=0$, but whose square does not $\int_{\mathbf{x}}^{\mathbf{x}+\mathbf{a}}\mu(\mathbf{x})\mu(\mathbf{x})\neq 0$, and random disorder with $\langle g(\mathbf{x})\rangle=0$, $\langle g(\mathbf{x})g(\mathbf{x^{\prime}})\rangle=g^{2}\delta(\mathbf{x}-\mathbf{x^{\prime}})$. Then by the Harris criterion harris_effect_1974, generic Gaussian random field disorder introduced through $\int\!\text{d}\vec{x}\text{d}t\,g(\vec{x}){\cal O}(\vec{x})$ is relevant if the scaling dimension of $\cal O$} is $\Delta_{\cal O}<\frac{1}{2}(d+1)$}. By analogy, for a chemical potential modulation around a charge neutral lattice coupling to ${\cal O}(\vec{x})=n(\vec{x})$ with dimension $[n]=d$, the modulation saturates this bound for $d=1$, but violates it for $d\geq 2$. Such a lattice modulation is thus marginal in $d=1$ and irrelevant in higher dimensions. [^4] As can be seen in Figure 2, the lattice modulation is marginally irrelevant for 1D lattices. The behavior of the charge density is similar as for 2D lattices. This is supported by an analysis of the near-horizon geometry that determines the DC response. Similar to 1D lattices, for $T=0$ the metric remains purely AdS4 in the IR. However, for a square lattice as we use, this is an isotropic AdS4 and there is no enhancement expected of the quantum critical conductivity. Indeed our numerics show that the 2D DC electrical conductivity approaches $\sigma_{\infty}$ as temperature decreases, though we do note that we are numerically limited by our numerics to go to temperatures lower than $T/\delta\mu=0.01$ while keeping $\delta\mu/G$ reasonably large.
@@
At the same time, the arguments for a trivial conductivity $\sigma_{DC}=\sigma_{\infty}$ at high temperatures in the case of 1D must break down for 2D lattices, as is readily confirmed. In particular, from the holographic point of view, Eq. 29 now allows for spatial variations in the horizon velocity field. Thus, while the unit cell average of currents perpendicular to the electric field $E_{x}$ must vanish ($\bar{J}_{y}=0$), local variations of $J_{y}(\vec{x})$ need not be zero. [^5] Recall that $\bar{\mu}=0$ causes the overall global Hall conductivity ($\sigma_{xy}=0$) to vanish, but local variations are still allowed. In fact, at high temperatures it will in general fluctuate, whereas at low temperatures we find $J_{y}(\vec{x})\to 0$. We show in the inset in Figure 6(a) that the variance of $J_{y}(\vec{x})$ indeed follows the same general shape as the electrical conductivity. We may interpret this variance in $J_{y}(\vec{x})$ as the consequence of the current following a curved path through the potential landscape, in such a way that the conductivity is increased. Though a holographic computation in the strong coupling regime cannot reveal this, the natural mechanism prompting such perpendicular variations in a weakly coupled (Boltzmann transport) system that grow with temperature would be thermally assisted fluctuations or thermally assisted local scattering.
@@
This picture is corroborated by a direct inspection of the current flows. Figure 7 shows both the unrenormalized lattice as well as the local electrical and thermal currents $J_{i}(\vec{x})$ and $Q_{i}(\vec{x})$ on the horizon, representing the quantities solved for in the Stokes equations 12.666Note that the extraction of the physical currents in the dual quantum critical theory from the horizon currents is based on the fact that the unit cell averages $\bar{J}_{i},\bar{Q}_{i}$ do not renormalize. The horizon values of the local expressions $J_{i}(\vec{x})$ and $Q_{i}(\vec{x})$ do renormalize and are in general not identical to their boundary counterparts. Nevertheless, the Stokes equations describing the problem give insight into the underlying physics. Indeed, we find a similar pattern in the boundary currents $J_{i}(\vec{x})$ and $Q_{i}(\vec{x})$ as computed by a (more numerically complicated) AC calculation for $\omega\to 0$. From this, an interesting picture emerges, which on physical grounds must have its origins in that the collision-less to hydrodynamic crossover in quantum critical systems can happen at larger scales than one would expect: it seems that each patch in the transverse plane can be thought of as a local (electric) Reissner-Nordström geometry, i.e., locally it is a collective quantum critical system at finite density, even though the global charge density vanishes. If so, one can consider the hydrodynamic results for the conductivities 25, assuming the expressions to be valid at some (suitably defined) local point $\vec{x}$, which should hold in an adiabatic approximation where the spatial variation is on a much larger scale than the onset of hydrodynamics: [^7] As the system is in thermal equilibrium, the temperature $T$ will still be $\vec{x}$-independent.
@@
Since for a finite density system, the electrical Drude weight is given by the (local) density squared, the local conductivity $\sigma_{DC}(\vec{x})$ is insensitive to the sign changes of the local chemical potential. This can be seen in Figure 7(a): the electrical current density $J_{i}(\vec{x})$ concentrates in regions of large $\left|\mu(\vec{x})\right|$ independent of the sign, and evades regions where $\mu(\vec{x})\approx 0$ (where the local conductivity is $\sigma=\sigma_{\infty}=1$). This is reminiscent of Effective Medium Theory, where the current takes the path of least resistance through a material with regions of locally strongly varying conductivity stroud_generalized_1975.888Note that this means that locally, transport is governed by a Drude-like pole, but there is no global hydrodynamics where a Drude mechanism can explain transport.
This same evading flow patterns are present in local probes of the thermal horizon current, albeit with an important twist. Figure 7(b) shows the thermal current $Q_{i}(\vec{x})$ when applying a constant thermal gradient in the $x$-direction. Keeping with our picture of local Reissner-Nordström geometries and local hydrodynamic transport, the thermal conductivity $\bar{\kappa}(\vec{x})$ will be proportional to $\frac{s(\vec{x})^{2}T}{{s(\vec{x})T}+{\mu(\vec{x})n(\vec{x})}}$ upon approximating $\epsilon+P=sT+\mu n$, cf. 32. Similar to the electrical conductivity, the thermal conductivity is therefore insensitive to the local sign of the chemical potential: however, now regions of large $\mu(\vec{x})$ are regions of weak transport and vice versa. Hence one has a weaker thermal current at large local $\left|\mu(\vec{x})\right|$, and a stronger current in regions of small local $\left|\mu(\vec{x})\right|$, in contrast to the electrical conductivity. [^9] The thermoelectric conductivity $\alpha$ — not shown here for brevity — is on the other hand sensitive to the sign of the local chemical potential and density (cf. Eq.25) and indeed, we find that the local current $J_{i}(\vec{x})$ when applying a thermal gradient (or thermal current $Q_{i}(\vec{x})$ for an external electric field) flows in a direction set by $\text{sgn}(\mu(\vec{x}))$, leading to $\alpha=\bar{\alpha}=0$ upon averaging (since $\bar{\mu}=0$).
@@ 5.2 AC transport in 2D charge-neutral lattices
Figure 8: AC transport in a 2D charge neutral lattice, $\mu(\vec{x})=\frac{1}{2}\delta\mu\left(\cos(Gx)+\cos(Gy)\right)$. (a-b) In contrast to the 1D lattice, the 2D AC electrical conductivity $\sigma(\omega)$ exhibits a weak zero-frequency peak at large lattices, in addition to the the Umklapp sound peak at $\omega\approx G/\sqrt{2}$ (gray dashed line), which now flattens out at larger temperatures. Inset: location of the new zero-frequency diffusive pole $\omega^{*}_{\text{elec}}$ along the imaginary axis. The solid gray line is a fit on the form $\operatorname{Im}\omega^{*}=a+b\,\delta\mu^{2}$ over the interval $\delta\mu/G\in(0,4]$. (c-d) As for 1D lattices, the 2D AC thermal conductivity $\bar{\kappa}(\omega)$ shows a similar mid-IR peak at larger lattice strengths arising from the collision of the Drude pole with the Umklapp charge diffusion pole. There is also an additional novel zero-frequency peak as for the electrical conductivity. Inset: location of the new zero-frequency diffusive pole $\omega^{*}_{\text{therm}}$ along the imaginary axis. The gray dashed line is a guide to the eye. The scaling with $\delta\mu$ is different from the corresponding pole in the electrical conductivity, and is thus of a different origin.
@@
With our understanding of the DC conductivities through an Effective Medium flow through channels of least resistance, we can attempt a qualitative explanation of the underlying physics. The essence of this Effective Medium type point of view is that one essentially has a composite of independent materials — in this case particle- and hole-filled regions created by the periodic chemical potential. Each independent region has a single fluid hydrodynamic description, meaning that there are effectively as many Drude modes as there are local fluid regions that weakly interact with each other. This viewpoint is supported by similar findings in charge-neutral graphene, where in a periodic background one has new Dirac points that are remnants of the local Fermi surface in each pocket parkNewGenerationMassless2008; chesler_conformal_2014. In a collective interacting system, the same should happen for the emergent hydrodynamics provided the lattice period is large enough and the boundaries sharp enough. In a 1D charge neutral periodic lattice these modes act in unison, however; most of the local contributions either add coherently or cancel each other perfectly and a single global hydrodynamics survives. In particular, there is no net charge flow leaving only a quantum critical electrical conductivity. In 2D, however, thanks to variance of the flow allowed by the perpendicular direction, this cancellation of each of the local hydrodynamic charge flow is no longer exact and a remnant Drude-like mode $\omega^{\ast}_{\text{elec}}$ survives. This mode exists in addition to the global Drude mode, which happens to collide with the global Umklapp charge diffusion mode. This “existence of more modes than conserved quantities” would be the hydrodynamic counterpart of the violation of the RG $c$-theorem in charge-neutral graphene in a periodic background. Naively, $c$ counts the number of massless Dirac modes, which in charge-neutral graphene in a periodic potential increases due to the emergence of the new Dirac points. This violation is allowed because the $c$-theorem relies on translational invariance, which is no longer present.
@@
On top of this novel emergent Drude-like mode of mesoscopic Effective Medium origin, we do recognize the same characteristics underlying 2D charge neutral quantum critical transport as in 1D. In the AC electrical conductivity $\sigma(\omega)$ at intermediate frequencies we see the Umklapped sound mode at $\omega=G/\sqrt{2}$. In contrast to 1D lattices, at temperatures $T\gg G$ the Umklapped sound mode now disappears. One possible explanation is that though the lattice is irrelevant and therefore more important at higher $T$, the coupling that allows sound from the energy channel to communicate with electrical transport does weaken at higher TT. [^11] It appears that the Umklapp sound pole moves to the upper half of the imaginary plane for $T\gtrsim 2G$, signaling a dynamical lattice instability. Similar such sort local chemical potential lattice type instabilities are known to occur in Bose-Einstein condensates in optical lattices (known as dynamical or modulational instabilities as opposed to thermodynamic instabilities), see e.g., sarloUnstableRegimesBoseEinstein2005; tasgalSoundWavesModulational2015. We have not investigated this further.
@@ 5.3 DC Magnetotransport in 2D charge-neutral lattices
Holding all other parameters ($T/G,\delta\mu/G)$ fixed, the most striking result is an almost linear dependence of the longitudinal magneto-resistance $\rho_{xx}=\frac{1}{\sigma_{xx}}$ on the magnetic field $B/T^{2}$ beyond $B\sim T^{2}$ independent of the size of $\delta\mu/G$: in Figure 9, we plot the logarithmic derivative of $\rho(B)$ to extract the local scaling exponent $\beta$ for $\rho_{xx}\sim\left|B\right|^{\beta}$. In the “Effective Medium regime”, $T\gtrsim G$, we obtain roughly linear dependence on the magnetic field. As we increase the temperature further, we find that the scaling is slightly superlinear with $\beta\simeq 1.25$ for most finite lattice amplitudes $\delta\mu/G>0$. Recalling that parity implies that the longitudinal resistivity can only be a function of $B^{2}$, the explanation might therefore be a quadrature-type relation $\rho_{xx}\sim({\rho_{xx}(B=0)^{2/\beta}+\gamma\mu_{B}B^{2}})^{\beta/2}$} well-known to occur in Effective Medium Theory parish_non-saturating_2003; hayes_scaling_2016; ramakrishnan_equivalence_2017; boyd_single-parameter_2019; ayres_incoherent_2021. Notably, we obtain emergent $B$-linear scaling without an ad hoc addition of mesoscopic disorder on top of our microscopic model. As we argued in the previous sector, the chemical potential modulation in a charge neutral background suffices to cause a flow to a ground state that consists of effectively independent particle-filled and hole-filled regions.

### v2

@@ 5 Thermoelectric transport in 2D charge-neutral lattices
We will show that there is a qualitatively novel contribution to the phenomenology of quantum critical transport in 2D charge-neutral lattices. In addition, the electrical conductivity has a modified RG-scaling from 1D. At the same time, the novel pole-collision physics behind 1D thermal transport does appear to continue to hold in 2D.
@@ 5.1 DC transport in 2D charge-neutral lattices
The difference in electrical transport is immediately evident in the DC conductivity $\sigma_{DC}$, as shown in Figure 6 (panel (a)): the electrical conductivity now increases with temperature instead of decreasing monotonically back down to the translationally invariant quantum critical value $\sigma_{\infty}$. This remarkable enhancement of the conductivity at higher temperatures can be understood if the lattice modulation acts as an irrelevant deformation. For a spatially dependent operator this is not straightforward to determine, but there is heuristic argument that suggests that this is indeed so. One can make an analogy between a periodic source of translational symmetry breaking which vanishes averaged over a lattice cell $\int_{\vec{x}}^{\vec{x}+\vec{a}}\mu(\vec{x})=0$ but whose square does not $\int_{\vec{x}}^{\vec{x}+\vec{a}}\mu(\vec{x})\mu(\vec{x})\neq 0$, and random disorder with $\langle g(\vec{x})\rangle=0$, $\langle g(\vec{x})g(\vec{x}^{\prime})\rangle=g^{2}\delta(\vec{x}-\vec{x}^{\prime})$. For such generic Gaussian random field disorder introduced through $\int\!\text{d}\vec{x}\text{d}t\,g(\vec{x}){\cal O}(\vec{x})$, the Harris criterion Harris (1974) reveals that it is relevant if the scaling dimension of $\cal O$} is $\Delta_{\cal O}<\frac{1}{2}(d+1)$}. For a randomly disordered chemical potential modulation around a charge neutral lattice coupling to ${\cal O}(\vec{x})=n(\vec{x})$ with dimension $[n]=d$, the modulation saturates this bound for $d=1$, but violates it for $d\geq 2$. Such chemical potential disorder is thus marginal in $d=1$ and irrelevant in higher dimensions. Although we here consider lattice modulations around zero average rather than random disorder, our numerics support this same division, suggesting that the vanishing of averages makes the lattice modulation act like disorder and that this is a good hint towards the underlying physics. [^6] As can be seen in Figure 2, the lattice modulation is marginally irrelevant for 1D lattices. The charge-density profile exhibits the same qualitative behavior for 2D lattices.
This is further supported by an analysis of the near-horizon geometry that determines the DC response. Similar to 1D lattices, for $T=0$ the metric remains purely AdS4 in the IR. However, for the square lattice we use, this is an isotropic AdS4 and there is no enhancement expected of the quantum critical conductivity. Indeed our numerics show that the 2D DC electrical conductivity approaches $\sigma_{\infty}$ as temperature decreases. This must be from a higher value $\sigma_{T}>\sigma_{\infty}$ by the bound Grozdanov et al. (2015), thus following an irrelevant RG flow. We do note that we are numerically limited by our numerics to go to temperatures lower than $T/\delta\mu=0.01$ while keeping $\delta\mu/G$ reasonably large.
At the same time, the arguments for a trivial conductivity $\sigma_{DC}=\sigma_{\infty}$ at high temperatures in the case of 1D must break down for 2D lattices, as is readily confirmed. In particular, from the holographic point of view, 29 now allows for spatial variations in the horizon velocity field. The Stokes equations 12 reduce in this limit to
$v^{i}\partial_{i}\rho^{H}+\partial_{i}(E^{i}+\partial^{i}w)$ $=0\,$, (31)
$\partial^{2}v_{i}+4\pi T\zeta_{i}-\partial_{i}p+\rho^{H}(E^{i}+\partial^{i}w)$ $=0\,$, (32)
@@
where we have used that all transport coefficients but $\rho^{H}$ become trivial and that $\partial_{i}v^{i}=0$. Thus, while the unit cell average of currents perpendicular to the electric field $E_{x}$ must vanish ($\bar{J}_{y}=0$), local variations of $J_{y}(\vec{x})$ need not be zero. [^7] Recall that $\bar{\mu}=0$ causes the global Hall conductivity ($\sigma_{xy}=0$) to vanish, but local variations are still allowed. In fact, at high temperatures it will in general fluctuate, whereas at low temperatures we find $J_{y}(\vec{x})\to 0$. We show in the inset in Figure 6(a) that the variance of $J_{y}(\vec{x})$ indeed follows the same general shape as the electrical conductivity. We may interpret this variance in $J_{y}(\vec{x})$ as the consequence of the current following a curved path through the potential landscape, in such a way that the conductivity is increased. Though a holographic computation in the strong coupling regime cannot reveal this, the natural mechanism prompting such perpendicular variations in a weakly coupled (Boltzmann transport) system that grow with temperature would be thermally assisted fluctuations or thermally assisted local scattering.
@@
This picture is corroborated by a direct inspection of the current flows. Figure 7 shows both the unrenormalized lattice as well as the local electrical and thermal currents $J_{i}(\vec{x})$ and $Q_{i}(\vec{x})$ on the horizon, representing the quantities solved for in the Stokes equations 12.888Note that the extraction of the physical currents in the dual quantum critical theory from the horizon currents is based on the fact that the unit cell averages $\bar{J}_{i},\bar{Q}_{i}$ do not renormalize. The horizon values of the local expressions $J_{i}(\vec{x})$ and $Q_{i}(\vec{x})$ do renormalize and are in general not identical to their boundary counterparts. Nevertheless, the Stokes equations describing the problem give insight into the underlying physics. We do in fact find a similar pattern in the boundary currents $J_{i}(\vec{x})$ and $Q_{i}(\vec{x})$ as computed by a (more numerically complicated) AC calculation for $\omega\to 0$. From this, an interesting picture emerges, which on physical grounds must have its origins in that the collision-less to hydrodynamic crossover in quantum critical systems can happen at larger scales than one would expect: it seems that each patch in the transverse plane can be thought of as a local (electric) Reissner-Nordström geometry, i.e., locally it is a collective quantum critical system at finite density, even though the global charge density vanishes. If so, one can consider the hydrodynamic results for the conductivities 25, assuming the expressions to be valid at some (suitably defined) local point $\vec{x}$, which should hold in an adiabatic approximation where the spatial variation is on a much larger scale than the onset of hydrodynamics: [^9] As the system is in thermal equilibrium, the temperature $T$ will still be $\vec{x}$-independent.
@@
Since for a finite density system, the electrical Drude weight is given by the (local) density squared, the local conductivity $\sigma_{DC}(\vec{x})$ is insensitive to the sign changes of the local chemical potential. This can be seen in Figure 7(a): the electrical current density $J_{i}(\vec{x})$ concentrates in regions of large $\left|\mu(\vec{x})\right|$ independent of the sign, and evades regions where $\mu(\vec{x})\approx 0$ (where the local conductivity is $\sigma=\sigma_{\infty}=1$). This is reminiscent of Effective Medium Theory, where the current takes the path of least resistance through a material with regions of locally strongly varying conductivity Stroud (1975). [^10] Note that this means that locally, transport is governed by a Drude-like pole, but there is no global hydrodynamics where a Drude mechanism can explain transport.
This same evading flow patterns are present in local probes of the thermal horizon current, albeit with an important twist. Figure 7(b) shows the thermal current $Q_{i}(\vec{x})$ when applying a constant thermal gradient in the $x$-direction. Keeping with our picture of local Reissner-Nordström geometries and local hydrodynamic transport, the thermal conductivity $\bar{\kappa}(\vec{x})$ will be proportional to $\frac{s(\vec{x})^{2}T}{{s(\vec{x})T}+{\mu(\vec{x})n(\vec{x})}}$ upon approximating $\epsilon+P=sT+\mu n$, cf. 34. Similar to the electrical conductivity, the thermal conductivity is therefore insensitive to the local sign of the chemical potential: however, now regions of large $\mu(\vec{x})$ are regions of weak transport and vice versa. Hence one has a weaker thermal current at large local $\left|\mu(\vec{x})\right|$, and a stronger current in regions of small local $\left|\mu(\vec{x})\right|$, in contrast to the electrical conductivity. [^11] The thermoelectric conductivity $\alpha$ — not shown here for brevity — is on the other hand sensitive to the sign of the local chemical potential and density (cf. 25) and indeed, we find that the local current $J_{i}(\vec{x})$ when applying a thermal gradient (or thermal current $Q_{i}(\vec{x})$ for an external electric field) flows in a direction set by $\text{sgn}(\mu(\vec{x}))$, leading to $\alpha=\bar{\alpha}=0$ upon averaging (since $\bar{\mu}=0$).
@@ 5.2 AC transport in 2D charge-neutral lattices
Figure 8: AC transport in a 2D charge neutral lattice, $\mu(\vec{x})=\frac{1}{2}\delta\mu\left(\cos(Gx)+\cos(Gy)\right)$. (a-b) In contrast to the 1D lattice, the 2D AC electrical conductivity $\sigma(\omega)$ exhibits a weak zero-frequency peak at large lattices, in addition to the the Umklapp sound peak at $\omega\approx G/\sqrt{2}$ (gray dashed line). Inset: location of the new zero-frequency diffusive pole $\omega^{*}_{\text{elec}}$ along the imaginary axis. The solid gray line is a fit on the form $\operatorname{Im}\omega^{*}=a+b\,\delta\mu^{2}$ over the interval $\delta\mu/G\in(0,4]$. (c-d) As for 1D lattices, the 2D AC thermal conductivity $\bar{\kappa}(\omega)$ shows a similar mid-IR peak at larger lattice strengths arising from the collision of the Drude pole with the Umklapp charge diffusion pole. There is also an additional novel zero-frequency peak as for the electrical conductivity. Inset: location of the new zero-frequency diffusive pole $\omega^{*}_{\text{therm}}$ along the imaginary axis. The gray dashed line is a guide to the eye. The scaling with $\delta\mu$ is different from the corresponding pole in the electrical conductivity, and is thus of a different origin.
@@
With our understanding of the DC conductivities through an Effective Medium flow through channels of least resistance, we can attempt a qualitative explanation of the underlying physics. The essence of this Effective Medium type point of view is that one essentially has a composite of independent materials — in this case particle-filled and hole-filled regions created by the periodic chemical potential. Each independent region has a single fluid hydrodynamic description, meaning that there are effectively as many Drude modes as there are local fluid regions that weakly interact with each other. This viewpoint is supported by similar findings in charge-neutral graphene, where in a periodic background one has new Dirac points that are remnants of the local Fermi surface in each pocket Park et al. (2008b); Chesler et al. (2014). In a collective interacting system, the same should happen for the emergent hydrodynamics provided the lattice period is large enough and the boundaries sharp enough. In a 1D charge neutral periodic lattice these modes act in unison, however; most of the local contributions either add coherently or cancel each other perfectly and a single global hydrodynamic description survives. In particular, there is no net charge flow leaving only a quantum critical electrical conductivity. In 2D, however, thanks to variance of the flow allowed by the perpendicular direction, this cancellation of each of the local hydrodynamic charge flow is no longer exact and a remnant Drude-like mode $\omega^{\ast}_{\text{elec}}$ survives. This mode exists in addition to the global Drude mode, which happens to collide with the global Umklapp charge diffusion mode. This “existence of more modes than conserved quantities” would be the hydrodynamic counterpart of the violation of the RG $c$-theorem in charge-neutral graphene in a periodic background. Naively, $c$ counts the number of massless Dirac modes, which in charge-neutral graphene in a periodic potential increases due to the emergence of the new Dirac points. This violation is allowed because the $c$-theorem relies on translational invariance, which is no longer present.
@@
On top of this novel emergent Drude-like mode of mesoscopic Effective Medium origin, we do recognize the same characteristics underlying 2D charge neutral quantum critical transport as in 1D. In particular, in the AC electrical conductivity $\sigma(\omega)$ at intermediate frequencies, we still see the Umklapped sound mode at $\omega=G/\sqrt{2}$.
@@ 5.3 DC magnetotransport in 2D charge-neutral lattices
Holding all other parameters ($T/G,\delta\mu/G)$ fixed, the most striking result is an almost linear dependence of the longitudinal magneto-resistance $\rho_{xx}=\frac{1}{\sigma_{xx}}$ on the magnetic field $B/T^{2}$ beyond $B\sim T^{2}$ independent of the size of $\delta\mu/G$: in Figure 9, we plot the logarithmic derivative of $\rho(B)$ to extract the local scaling exponent $\beta$ for $\rho_{xx}\sim\left|B\right|^{\beta}$. In the “Effective Medium regime”, $T\gtrsim G$, we obtain roughly linear dependence on the magnetic field. As we increase the temperature further, we find that the scaling stays quasi-linear over a larger regime. Recalling that parity implies that the longitudinal resistivity can only be a function of $B^{2}$, the explanation might therefore be a quadrature-type relation $\rho_{xx}\sim({\rho_{xx}(B=0)^{2/\beta}+\gamma\mu_{B}B^{2}})^{\beta/2}$} well-known to occur in Effective Medium Theory Parish and Littlewood (2003); Hayes et al. (2016); Ramakrishnan et al. (2017); Boyd and Phillips (2019); Ayres et al. (2021). Notably, we obtain emergent $B$-linear scaling without an ad hoc addition of mesoscopic disorder on top of our microscopic model. As we argued in the previous sector, the chemical potential modulation in a charge neutral background suffices to cause a flow to a ground state that consists of effectively independent particle-filled and hole-filled regions.